OpenStudy (anonymous):

Help (ln(2x))(3ln(2x)-4)=5

OpenStudy (anonymous):

let)ln(2x) = t , then t(3t-4)=5 i) 3t^2 - 4t - 5 = 0 (x>0) then you will see that thi eq cannot factor. so use quadratic formula explicity t= (\[\frac{ 4\pm \sqrt{19} }{ 3 }\] so, t=ln2x actually. answer => x=[e^{4+root(19)}/3] or x=[e^{4-root(19)}/3]